# Approximation of magnetic Schrödinger operators with $\delta$-interactions supported on networks

### v1

@@ Abstract.
This paper deals with the approximation of a magnetic Schrödinger operator with a singular $\delta$-potential that is formally given by $(i\nabla+A)^{2}+Q+\alpha\delta_{\Sigma}$ by Schrödinger operators with regular potentials in the norm resolvent sense. This is done for $\Sigma$ being the finite union of $C^{2}$-hypersurfaces, for coefficients $A$, $Q$, and $\alpha$ under minimal assumptions such that the associated quadratic forms are closed and sectorial, and $Q$ and $\alpha$ are allowed to be complex-valued functions. In particular, $\Sigma$ can be a graph in $\mathbb{R}^{2}$ or the boundary of a piecewise $C^{2}$-domain. Moreover, spectral implications of the mentioned convergence result are discussed.
@@ 1. Introduction
Schrödinger operators with singular $\delta$-potentials supported on sets of measure zero play an important role in mathematical physics. Singular potentials supported on points were already used in the early days of quantum mechanics [29, 39] and later systematically analyzed from a mathematical point of view, see the monograph [1] and the references therein. In the last three decades, there was a substantial interest in singular potentials supported on hypersurfaces $\Sigma$ in $\mathbb{R}^{n}$, $n\geq 2$, like curves in $\mathbb{R}^{2}$ or surfaces in $\mathbb{R}^{3}$, see the review paper [16], the monograph [18], or the research papers [5, 6, 9, 14, 15, 17, 21, 33], to mention just a few selected works. Such operators are formally given by
@@
and they describe for attractive interactions (i.e. $\alpha<0$) the propagation of a quantum particle on $\Sigma$ under the influence of a magnetic vector potential $A$ and an electric potential $Q$ allowing the tunneling of the particle between different parts of $\Sigma$, see [5, 16]; from this point of view, graph like structures $\Sigma\subset\mathbb{R}^{2}$ are of particular interest. Moreover, operators of the form $H_{0,0,\alpha}$ are useful in the design of photonic crystals with prescribed spectral gaps [22, 26].
@@
Of course, the operators with the singular potentials are only idealized replacements of more realistic operators with regular potentials, and so the relation of the regular and the singular models was justified via an approximation procedure already in [29]. For $A=Q=0$ the approximation of one, two, and three-dimensional Schrödinger operators with $\delta$-point interactions by Schrödinger operators with scaled regular potentials in the norm resolvent sense was treated systematically in the 1970s and 1980s, see the monograph [1] and the references therein; the latter convergence implies, in particular, the convergence of the spectra and of functions of the involved operators [27, 40] and thus, it justifies the usage of the idealized models as good approximations of their regular counterparts. The approximation of Schrödinger operators with $\delta$-potentials supported on the sphere in $\mathbb{R}^{3}$ was considered in [2, 37], then the situation that $\Sigma$ is the boundary of a star shaped domain and $\alpha$ is real-valued and smooth was investigated in [36], and unbounded and $C^{2}$-smooth $\Sigma$ in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ and constant $\alpha\in\mathbb{R}$ were treated in [17, 21]. It is worth mentioning that in [38] Laplacians with $\delta$- (and other measure valued) potentials supported on quite general geometries were approximated, but only in the strong resolvent sense, which does not imply the convergence of the spectra of the approximating operators in the Hausdorff metric. Finally, in [4] the approximation of the Schrödinger operator with a $\delta$-potential of real-valued strength $\alpha\in L^{\infty}(\Sigma)$ on one bounded or unbounded $C^{2}$-smooth hypersurface $\Sigma\subset\mathbb{R}^{n}$, $n\geq 2$, was shown, and in [5] the approximation of a two-dimensional Schrödinger operator with a homogeneous magnetic field and an electric $\delta$-potential of real-valued strength $\alpha\in L^{\infty}(\Sigma)$ supported on a bounded $C^{1,1}$-smooth curve was treated.
@@
The main motivation in this paper is to show that the operator $H_{A,Q,\alpha}$ formally given by (1.1) can be approximated in the norm resolvent sense by Schrödinger operators with scaled regular potentials under minimal assumptions on $A,Q$, and $\alpha$, and for a wide class of interaction supports $\Sigma$. Concerning the coefficients, of course from the point of view of quantum mechanics, real-valued functions that are associated with self-adjoint operators are favored, but the analysis in this paper works equally well for complex-valued functions $Q$ and $\alpha$, yielding a relation to works on Schrödinger operators with non-self-adjoint $\delta$-potentials [10, 24, 28]. Moreover, the interaction support $\Sigma$ will be allowed to be the finite union of bounded or unbounded $C^{2}$-smooth hypersurfaces, and so classes of graph-like structures in $\mathbb{R}^{2}$, that are particularly important for the application to leaky quantum graphs, and piecewise $C^{2}$-smooth $\Sigma\subset\mathbb{R}^{n}$, which are an important subclass of Lipschitz domains, are contained.
@@
To set the stage in a rigorous way, let $N,n\in\mathbb{N}$ be fixed with $n\geq 2$ and $\Sigma^{(1)},\dots,\Sigma^{(N)}\subset\mathbb{R}^{n}$ be hypersurfaces as in Definition 2.5 with normal vector fields $\nu^{(1)},\dots,\nu^{(N)}$ such that the Hausdorff measure of $\Sigma^{(k)}\cap\Sigma^{(l)}$ is zero for $k\neq l$. Then, the network $\Sigma$ is
@@
This space is studied in detail in Section 2.1, where it is shown that $\mathcal{H}_{A,Q}$, when endowed with a proper norm, is a Hilbert space and that it is indeed independent of $Q_{2}$. Moreover, in Section 2.2 it is proved that functions $f\in\mathcal{H}_{A,Q}$ posses Dirichlet traces $f|_{\Sigma^{(k)}}\in L^{q}(\Sigma^{(k)})$ with $q=2\frac{n-1}{n-2+\mu}$ and $\mu\in(0,1]$. Eventually, assume that
@@
is injective, see Proposition 2.6. Introduce for $\varepsilon\in(0,\beta]$ the set $\Omega_{\varepsilon}^{(k)}:=\iota^{(k)}(\Sigma^{(k)}\times(-\varepsilon,\varepsilon))$, and choose a function $V^{(k)}\in L^{p}(\mathbb{R}^{n})+L^{\infty}(\mathbb{R}^{n})$ with $p=\frac{n-1}{1-\gamma}$ for a $\gamma\in(0,1)$ that is supported in $\Omega_{\beta}^{(k)}$. Then, one can define for $k\in\{1,\dots,N\}$
@@
$\alpha^{(k)}(x_{\Sigma^{(k)}}):=\int_{-\beta}^{\beta}V^{(k)}(x_{\Sigma^{(k)}}+t\nu^{(k)}(x_{\Sigma^{(k)}}))\textup{d}\sigma(x_{\Sigma^{(k)}}).$ (1.11)
@@
with a constant $C=C(\lambda,\varepsilon_{0})>0$ is true. In particular, $H_{A,Q,V_{\varepsilon}}$ converges in the norm resolvent sense to $H_{A,Q,\alpha}$.
The proof of Theorem 1.1 follows ideas that were used in [5, Appendix B] and relies on an estimate of $\mathfrak{h}_{A,Q,V_{\varepsilon}}-\mathfrak{h}_{A,Q,\alpha}$ in a suitable topology and on applying an abstract result from [27, Section VI.3]. In particular, the fact that only estimates of the associated forms are used allows to treat non-smooth coefficients $A,Q,\alpha$ and networks $\Sigma$ and so, Theorem 1.1 improves the results on the approximation of Schrödinger operators with $\delta$-potentials supported on hypersurfaces in several directions:
@@
    The interaction strengths are allowed to be non-real and belong to $L^{p}$-spaces with optimal $p$ so that the operator $H_{A,Q,\alpha}$ can be defined with the help of its associated quadratic form.
@@
    Additive electric and magnetic potentials $A$ and $Q$, that are independent of $\varepsilon$, are allowed, and these potentials belong to optimal classes of potentials such that the unperturbed operators $H_{A,Q,0}$ can be defined as $m$-sectorial operators with the help of their associated form.
It is worth to note that also for the particularly important self-adjoint case Theorem 1.1 improves the known approximation results in all points (a)–(c) listed above.
@@
with a constant $C=C(\lambda,\varepsilon_{0})>0$ is true. In particular, $H_{A,Q,V_{\varepsilon}}$ converges in the norm resolvent sense to $H_{A,Q,\alpha}$.
@@
In this paper, $C$ denotes a generic constant that may change its value in-between lines. For an open set $\Omega\subset\mathbb{R}^{n}$ the set $H^{1}(\Omega)$ is the $L^{2}$-based Sobolev space of once weakly differentiable functions and $C^{k}_{b}(\Omega)$, $k\in\mathbb{N}$, is the set of all $k$ times continuously differentiable functions with bounded partial derivatives up to order $k$. For a closed linear operator $A$ in a Hilbert space, its domain, resolvent set, spectrum, and point spectrum are denoted by $\mathrm{dom}\,A$, $\rho(A)$, $\sigma(A)$, and $\sigma_{\textup{p}}(A)$. If $A$ is self-adjoint, then its discrete and essential spectrum are $\sigma_{\textup{disc}}(A)$ and $\sigma_{\textup{ess}}(A)$, respectively.
Acknowledgement.
The author thanks Jussi Behrndt and Petr Siegl for fruitful discussions. This research was funded in part by the Austrian Science Fund (FWF) 10.55776/P 33568-N. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@ 2.1. The space $\mathcal{H}_{A,Q}$
$\begin{split}\bigl{|}\mathfrak{q}_{2}[f]\bigr{|}&\leq\big{\|}|Q_{2}|^{1/2}|f|\big{\|}_{L^{2}(\mathbb{R}^{n})}^{2}\\ &\leq\delta\big{\|}\nabla|f|\big{\|}_{L^{2}(\mathbb{R}^{n})}^{2}+(C_{\delta}+\delta)\big{\|}|f|\big{\|}_{L^{2}(\mathbb{R}^{n})}^{2}\\ &\leq\delta\|f\|_{\mathcal{H}_{A,Q}}^{2}+(C_{\delta}+\delta)\big{\|}f\big{\|}_{L^{2}(\mathbb{R}^{n})}^{2}.\end{split}$
@@
The only property of the term $Q_{2}$ that we need in this paper and, in particular, to prove Theorem 1.1, is the one from Lemma 2.3. As seen in the proof, the statement remains true, if $Q_{2}$ is such that the form $\mathfrak{q}_{2}$ in (2.5) is form bounded with respect to the form $\mathcal{H}_{A,Q}\ni f\mapsto\|f\|_{\mathcal{H}_{A,Q}}^{2}$ with form bound smaller than one. Hence, Theorem 1.1 remains true for all $Q_{2}$ which have the latter property.
@@ 2.2. Description of hypersurfaces and their tubular neighborhoods
In this section the hypersurfaces that build the network being the support of the $\delta$-potential are described, and then a suitable variant of the trace theorem for these hypersurfaces and the space $\mathcal{H}_{A,Q}$ from (2.1) is discussed. Finally, some properties of the map $\iota$ in (1.8) which describes the tubular neighborhood of the network are collected. Since for the proof of Theorem 1.1 similar ideas as in [8, Section 3.2] are used, we follow closely the definition of hypersurfaces from [8]:
@@
We say that $\Sigma\subset\mathbb{R}^{n}$ is an admissible $C^{2}$-hypersurface, if there exist $p\in\mathbb{N}$, open sets $\Omega,V_{1},\dots,V_{p},W_{1},\dots,W_{p}\subset\mathbb{R}^{n}$, mappings $\zeta_{1},\dots,\zeta_{p}\in C^{2}_{b}(\mathbb{R}^{n-1};\mathbb{R})$, rotation matrices $\kappa_{1},\dots,\kappa_{p}\in\mathbb{R}^{n\times n}$, and $\varepsilon_{1}>0$ such that the following holds:
@@
    $W_{k}\cap\Omega=W_{k}\cap\{\kappa_{k}(x^{\prime},x_{n}):x^{\prime}\in\mathbb{R}^{n-1},x_{n}<\zeta_{k}(x^{\prime})\}$ for all $k\in\{1,\dots,p\}$.
@@
Roughly speaking, conditions (i)–(iii) in the above definition mean that the set $\Omega$ is of the same form as the sets $\Omega_{+}$ considered in [8], while conditions (iv)&(v) mean that the admissible hypersurface $\Sigma$ is a subset of $\partial\Omega$ with good properties, so that an integral over $\Sigma$ is well-defined. In particular, (relatively open subsets of) $C^{2}_{b}$-hypographs and the boundaries of compact $C^{2}$-domains are admissible $C^{2}$-hypersurfaces.
@@
An important role will be played by the Weingarten map or shape operator $W$ on an admissible $C^{2}$-hypersurface $\Sigma$. This is the linear map at the tangent space $T_{x_{\Sigma}}$ for $x_{\Sigma}=\kappa_{k}(x^{\prime},\zeta_{k}(x^{\prime}))\in\Sigma$ via its action on a basis element $\partial_{j}\kappa_{k}(x^{\prime},\zeta_{k}(x^{\prime}))$ of $T_{x_{\Sigma}}$,
$W(x_{\Sigma})\partial_{j}\kappa_{k}(x^{\prime},\zeta_{k}(x^{\prime}))=-\partial_{j}\nu\bigl{(}\kappa_{k}(x^{\prime},\zeta_{k}(x^{\prime}))\bigr{)},$ (2.7)
@@
Let $\Sigma\subset\mathbb{R}^{n}$ be an admissible $C^{2}$-hypersurface and $q=2\frac{n-1}{n-2+\mu}$ for $\mu\in(0,1]$. Then, the map $C_{0}^{\infty}(\mathbb{R}^{n})\ni f\mapsto f|_{\Sigma}$ admits a unique continuous extension $\gamma_{D}:\mathcal{H}_{A,Q}\rightarrow L^{q}(\Sigma)$ and for all $\delta>0$ there exists $C_{\delta}>0$ such that
@@
$\begin{split}\|f|_{\Sigma}\|_{L^{q}(\Sigma)}^{2}&=\bigl{\|}|f|\big{|}_{\Sigma}\bigr{\|}_{L^{q}(\Sigma)}^{2}\\ &\leq C\bigl{\|}|f|\big{|}_{\Sigma}\bigr{\|}_{H^{1/2-\mu/2}(\Sigma)}^{2}\\ &\leq C\bigl{\|}|f|\big{|}_{\Sigma}\bigr{\|}_{H^{1/2-\mu^{\prime}/2}(\Sigma)}^{2}\\ &\leq C\bigl{\|}|f|\bigr{\|}_{H^{1-\mu^{\prime}/2}(\mathbb{R}^{n})}^{2}\\ &\leq\delta\bigl{\|}|f|\bigr{\|}_{H^{1}(\mathbb{R}^{n})}^{2}+C_{\delta}\bigl{\|}|f|\bigr{\|}_{L^{2}(\mathbb{R}^{n})}^{2}\\ &\leq\delta\|f\|_{\mathcal{H}_{A,Q}}^{2}+C_{\delta}\|f\|_{L^{2}(\mathbb{R}^{n})}^{2},\end{split}$ (2.10)

### v2

@@ Abstract.
This paper deals with the approximation of a magnetic Schrödinger operator with a singular $\delta$-potential that is formally given by $(i\nabla+A)^{2}+Q+\alpha\delta_{\Sigma}$ by Schrödinger operators with regular potentials in the norm resolvent sense. This is done for $\Sigma$ being the finite union of $C^{2}$-hypersurfaces, for coefficients $A$, $Q$, and $\alpha$ under almost minimal assumptions such that the associated quadratic forms are closed and sectorial, and $Q$ and $\alpha$ are allowed to be complex-valued functions. In particular, $\Sigma$ can be a graph in $\mathbb{R}^{2}$ or the boundary of a piecewise $C^{2}$-domain. Moreover, spectral implications of the mentioned convergence result are discussed.
@@ 1. Introduction
Schrödinger operators with singular $\delta$-potentials supported on sets of measure zero play an important role in mathematical physics. Singular potentials supported on points were already used in the early days of quantum mechanics [29, 39] and later systematically analyzed from a mathematical point of view, see the monograph [1] and the references therein. In the last three decades, there was a substantial interest in singular potentials supported on hypersurfaces $\Sigma$ in $\mathbb{R}^{n}$, $n\geq 2$, such as curves in $\mathbb{R}^{2}$ or surfaces in $\mathbb{R}^{3}$, see the review paper [16], the monograph [19], or the research papers [5, 6, 9, 14, 15, 17, 18, 33], to mention just a few selected works. Such operators are formally given by
@@
and they describe for attractive interactions (i.e. for $\alpha<0$) the propagation of a quantum particle on $\Sigma$ under the influence of a magnetic vector potential $A$ and an electric potential $Q$ allowing the tunneling of the particle between different parts of $\Sigma$, see [5, 16]; from this point of view, graph like structures $\Sigma\subset\mathbb{R}^{2}$ are of particular interest. Moreover, operators of the form $H_{0,0,\alpha}$ are useful in the design of photonic crystals with prescribed spectral gaps [22, 26].
@@
Of course, the operators with the singular potentials are only idealized replacements of more realistic operators with regular potentials. The relation of the regular and the singular models was justified via an approximation procedure already in [29]. For $A=Q=0$ the approximation of one, two, and three-dimensional Schrödinger operators with $\delta$-point interactions by Schrödinger operators with scaled regular potentials in the norm resolvent sense was treated systematically in the 1970s and 1980s, see the monograph [1] and the references therein; the latter convergence implies, in particular, the convergence of the spectra and of functions of the involved operators [27, 40] and thus, it justifies the usage of the idealized models as good approximations of their regular counterparts. The approximation of Schrödinger operators with $\delta$-potentials supported on the sphere in $\mathbb{R}^{3}$ was considered in [2, 37], then the situation that $\Sigma$ is the boundary of a star shaped domain and $\alpha$ is real-valued and smooth was investigated in [36], and unbounded and $C^{2}$-smooth $\Sigma$ in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ and constant $\alpha\in\mathbb{R}$ were treated in [17, 18]. It is worth mentioning that in [38] Laplacians with $\delta$- (and other measure valued) potentials supported on quite general geometries were approximated, but only in the strong resolvent sense, which does not imply the convergence of the spectra of the approximating operators in the Hausdorff metric. Finally, in [4] the approximation of the Schrödinger operator with a $\delta$-potential of real-valued strength $\alpha\in L^{\infty}(\Sigma)$ on one bounded or unbounded $C^{2}$-smooth hypersurface $\Sigma\subset\mathbb{R}^{n}$, $n\geq 2$, was shown, and in [5] the approximation of a two-dimensional Schrödinger operator with a homogeneous magnetic field and an electric $\delta$-potential of real-valued strength $\alpha\in L^{\infty}(\Sigma)$ supported on a bounded $C^{1,1}$-smooth curve was treated; note that [5] is the only work known to the author where the approximation of $H_{A,Q,\alpha}$ for $A\neq 0$ was treated.
@@
The main motivation in this paper is to show that the operator $H_{A,Q,\alpha}$ formally given by (1.1) can be approximated in the norm resolvent sense by Schrödinger operators with scaled regular potentials under mild assumptions on $A,Q$, and $\alpha$, and for a wide class of interaction supports $\Sigma$. Concerning the coefficients, from the point of view of quantum mechanics real-valued functions that are associated with self-adjoint operators are favored, but the analysis in this paper works equally well for complex-valued functions $Q$ and $\alpha$, yielding a relation to works on Schrödinger operators with non-real $\delta$-potentials [10, 24, 28]. Moreover, the interaction support $\Sigma$ will be allowed to be the finite union of bounded or unbounded $C^{2}$-smooth hypersurfaces, and so classes of graph-like structures in $\mathbb{R}^{2}$, that are particularly important for the application to leaky quantum graphs, and piecewise $C^{2}$-smooth $\Sigma\subset\mathbb{R}^{n}$, which are an important subclass of Lipschitz domains, are contained.
@@
To set the stage in a rigorous way, let $N,n\in\mathbb{N}$ be fixed with $n\geq 2$ and $\Sigma^{(1)},\dots,\Sigma^{(N)}\subset\mathbb{R}^{n}$ be bounded or unbounded hypersurfaces as in Definition 2.5 with normal vector fields $\nu^{(1)},\dots,\nu^{(N)}$ such that the Hausdorff measure of $\Sigma^{(k)}\cap\Sigma^{(l)}$ is zero for $k\neq l$. Then, the network $\Sigma$ is
@@
This space is studied in detail in Section 2.1, where it is shown that $\mathcal{H}_{A,Q}$, when endowed with a proper norm, is a Hilbert space and that the domain of the quadratic form which is associated with the multiplication by $Q_{2}$ contains $\mathcal{H}_{A,Q}$; cf. Lemma 2.3. Moreover, in Section 2.2 it is proved that functions $f\in\mathcal{H}_{A,Q}$ possess Dirichlet traces $f|_{\Sigma^{(k)}}\in L^{q}(\Sigma^{(k)})$ with $q=2\frac{n-1}{n-2+\mu}$ and $\mu\in(0,1]$. Eventually, assume that
@@
is injective, see Proposition 2.7. Introduce for $\varepsilon\in(0,\beta]$ the set $\Omega_{\varepsilon}^{(k)}:=\iota^{(k)}(\Sigma^{(k)}\times(-\varepsilon,\varepsilon))$, and choose a function $V^{(k)}\in L^{p}(\mathbb{R}^{n})+L^{\infty}(\mathbb{R}^{n})$ with $p=\frac{n-1}{1-\gamma}$ for a $\gamma\in(0,1)$ that is supported in $\Omega_{\beta}^{(k)}$. Then, define for $k\in\{1,\dots,N\}$
@@
$\alpha^{(k)}(x_{\Sigma^{(k)}}):=\int_{-\beta}^{\beta}V^{(k)}(x_{\Sigma^{(k)}}+t\nu^{(k)}(x_{\Sigma^{(k)}}))\textup{d}t.$ (1.11)
@@
with a constant $C=C(\lambda,\varepsilon_{0})>0$ is true. In particular, $H_{A,Q,V_{\varepsilon}}$ converges in the norm resolvent sense to $H_{A,Q,\alpha}$, as $\varepsilon\searrow 0$.
The proof of Theorem 1.1 follows a strategy that was used in [5, Appendix B] and relies on an estimate of $\mathfrak{h}_{A,Q,V_{\varepsilon}}-\mathfrak{h}_{A,Q,\alpha}$ in a suitable topology and on applying an abstract result from [27, Section VI.3]. In particular, the fact that only estimates of the associated forms are used allows to treat non-smooth coefficients $A,Q,\alpha$ and networks $\Sigma$ and so, Theorem 1.1 improves the results on the approximation of Schrödinger operators with $\delta$-potentials supported on hypersurfaces in several directions:
@@
    The interaction strengths are allowed to be non-real and belong to $L^{p}$-spaces with almost optimal $p$ so that the operator $H_{A,Q,\alpha}$ can be defined with the help of its associated quadratic form.
@@
    Additive electric and magnetic potentials $Q$ and $A$, that are independent of $\varepsilon$, are allowed, and these potentials belong to optimal classes of potentials such that the unperturbed operator $H_{A,Q,0}$ can be defined as an $m$-sectorial operator with the help of the associated form.
It is worth to note that also for the particularly important self-adjoint case Theorem 1.1 improves the known approximation results in all points (a)–(c) listed above, as in the previous literature similar results were discussed only for one hypersurface $\Sigma$ (instead of a network) in [4, 17, 18, 36, 37], bounded interaction strengths $\alpha\in L^{\infty}(\Sigma)$ in [4, 17, 18, 36, 37], the homogeneous magnetic vector potential $A$ in dimension $d=2$ in [5], and bounded electric potentials $Q\in L^{\infty}(\mathbb{R}^{d})$ in [4].
@@
with a constant $C=C(\lambda,\varepsilon_{0})>0$ is true. In particular, $H_{A,Q,V_{\varepsilon}}$ converges in the norm resolvent sense to $H_{A,Q,\alpha}$, as $\varepsilon\searrow 0$.
@@
In this paper, $C$ denotes a generic constant that may change its value in between lines. For an open set $\Omega\subset\mathbb{R}^{n}$ the set $H^{1}(\Omega)$ is the $L^{2}$-based Sobolev space of once weakly differentiable functions and $C^{k}_{b}(\Omega)$, $k\in\mathbb{N}$, is the set of all $k$ times continuously differentiable functions with bounded partial derivatives up to order $k$. For a closed linear operator $A$ in a Hilbert space, its domain, resolvent set, spectrum, and point spectrum are denoted by $\mathrm{dom}\,A$, $\rho(A)$, $\sigma(A)$, and $\sigma_{\textup{p}}(A)$. If $A$ is self-adjoint, then its discrete and essential spectrum are $\sigma_{\textup{disc}}(A)$ and $\sigma_{\textup{ess}}(A)$, respectively.
@@ 2.1. The space $\mathcal{H}_{A,Q}$
$\begin{split}\bigl|\mathfrak{q}_{2}[f]\bigr|&\leq\big\||Q_{2}|^{1/2}|f|\big\|_{L^{2}(\mathbb{R}^{n})}^{2}\\ &\leq\delta\big\|\nabla|f|\big\|_{L^{2}(\mathbb{R}^{n})}^{2}+(C_{\delta}+\delta+1)\big\||f|\big\|_{L^{2}(\mathbb{R}^{n})}^{2}\\ &\leq\delta\|f\|_{\mathcal{H}_{A,Q}}^{2}+(C_{\delta}+\delta+1)\big\|f\big\|_{L^{2}(\mathbb{R}^{n})}^{2}.\end{split}$
@@
The only property of the term $Q_{2}$ that is needed in this paper and, in particular, to prove Theorem 1.1, is the one from Lemma 2.3. As seen in the proof, the statement remains true, if $Q_{2}$ is such that the form $\mathfrak{q}_{2}$ in (2.5) is form bounded with respect to the form $\mathcal{H}_{A,Q}\ni f\mapsto\|f\|_{\mathcal{H}_{A,Q}}^{2}$ with form bound smaller than one. Hence, Theorem 1.1 remains true for all $Q_{2}$ which have the latter property.
@@ 2.2. Description of hypersurfaces and their tubular neighborhoods
In this section the hypersurfaces that build the network being the support of the $\delta$-potential are described, and some properties of the map $\iota$ in (1.8) which describes the tubular neighborhood of the network are collected. Finally, a suitable variant of the trace theorem for these hypersurfaces and the space $\mathcal{H}_{A,Q}$ from (2.1) is discussed. Since for the proof of Theorem 1.1 similar ideas as in [8, Section 3.2] are used, we follow closely the definition of hypersurfaces from [8]:
@@
We say that $\Sigma\subset\mathbb{R}^{n}$ is an admissible $C^{2}$-hypersurface, if there exist a number $p\in\mathbb{N}$, open sets $\Omega,V_{1},\dots,V_{p},W_{1},\dots,W_{p}\subset\mathbb{R}^{n}$, real-valued mappings $\zeta_{1},\dots,\zeta_{p}\in C^{2}_{b}(\mathbb{R}^{n-1})$, rotation matrices $\kappa_{1},\dots,\kappa_{p}\in\mathbb{R}^{n\times n}$, and $\varepsilon_{1}>0$ such that the following holds:
@@
    $W_{k}\cap\Omega=W_{k}\cap\Omega_{k}$ for all $k\in\{1,\dots,p\}$, where the notation $\Omega_{k}=\{\kappa_{k}(x^{\prime},x_{n}):x^{\prime}\in\mathbb{R}^{n-1},x_{n}<\zeta_{k}(x^{\prime})\}$ is used.
@@
Note that the sets $\Omega_{1},\dots,\Omega_{p}$ appearing in item (iii) of the previous definition are rotated Lipschitz hypographs as in [35, Chapter 3], where the function $\zeta_{k}$ belongs to $C^{2}_{b}(\mathbb{R}^{n-1})$. Roughly speaking, conditions (i)–(iii) in the above definition mean that the set $\Omega$ is of the same form as the sets $\Omega_{+}$ considered in [8], while conditions (iv)&(v) mean that the admissible hypersurface $\Sigma$ is a subset of $\partial\Omega$ with good properties, so that an integral over $\Sigma$ is well-defined. In particular, (relatively open subsets of) $C^{2}_{b}$-hypographs and the boundaries of compact $C^{2}$-domains are admissible $C^{2}$-hypersurfaces. It will be important for our purposes that item (ii) in the previous definition implies in the case of unbounded hypersurfaces that different ends can not be too close to each other. A typical example of an unbounded curve in $\mathbb{R}^{2}$ which violates this condition is stated in the following example:
Example 2.6.
Let $\Omega\subset\mathbb{R}^{2}$ be an unbounded domain with boundary $\Sigma=\partial\Omega$ which can be described as in items (i) & (iii) of Definition 2.5 and which has the two ends $\pm(\sin^{2}(x)+x^{-1})$ for $x$ sufficiently large. Note that the two ends must be the boundaries of two different sets $\Omega_{k}$ as in Definition 2.5 (iii), as they can not be the image of one function $\mathbb{R}\ni t\mapsto\kappa(t,\zeta(t))$ with a rotation matrix $\kappa$ and $\zeta\in C^{2}_{b}(\mathbb{R})$; without loss of generality, the two ends are parts of the boundaries of
$\Omega_{1}:=\big\{(x,x_{2}):x\in\mathbb{R},x_{2}<\sin^{2}(x)+x^{-1}\big\}$
and
$\Omega_{2}:=\big\{(-x,-x_{2}):x\in\mathbb{R},x_{2}<\sin^{2}(x)-x^{-1}\big\}.$
Set $x_{n}=(n\pi,(n\pi)^{-1})$. Note that on the one hand item (iii) in Definition 2.5 implies for the associated sets $W_{1},W_{2}$ that $W_{k}\cap\Omega_{k}\subset\Omega$, $k\in\{1,2\}$. On the other hand, $x_{n}\in\partial\Omega_{1}$ for $n$ sufficiently large and for a fixed $\varepsilon_{1}>0$ one has $B(x_{n},\varepsilon_{1})\cap\Omega_{1}\not\subset\Omega$, implying $B(x_{n},\varepsilon_{1})\not\subset W_{1}$; in the same way, one argues that $B(x_{n},\varepsilon_{1})\not\subset W_{2}$. Therefore, this curve $\Sigma$ does not satisfy assertion (ii) in Definition 2.5.
@@
An important role will be played by the Weingarten map or shape operator $W$ on an admissible $C^{2}$-hypersurface $\Sigma$. This is the linear map at the tangent space $T_{x_{\Sigma}}$ for $x_{\Sigma}=\kappa_{k}(x^{\prime},\zeta_{k}(x^{\prime}))\in\Sigma$ via its action on a basis element $\partial_{x^{\prime}_{j}}\kappa_{k}(x^{\prime},\zeta_{k}(x^{\prime}))$ of $T_{x_{\Sigma}}$,
$W(x_{\Sigma})\partial_{x^{\prime}_{j}}\kappa_{k}(x^{\prime},\zeta_{k}(x^{\prime}))=-\partial_{x^{\prime}_{j}}\nu\bigl(\kappa_{k}(x^{\prime},\zeta_{k}(x^{\prime}))\bigr),\quad j\in\{1,\dots,n-1\},$ (2.7)
@@
Let $\Sigma\subset\mathbb{R}^{n}$ be an admissible $C^{2}$-hypersurface and $q=2\frac{n-1}{n-2+\mu}$ for $\mu\in(0,1]$. Then, the map $C_{0}^{\infty}(\mathbb{R}^{n})\ni f\mapsto f|_{\Sigma}$ admits a unique continuous extension $\gamma_{D}:\mathcal{H}_{A,Q}\rightarrow L^{q}(\Sigma)$ and for all $\delta>0$ there exists $C_{\delta}>1$ such that
@@
$\begin{split}\|f|_{\Sigma}\|_{L^{q}(\Sigma)}^{2}&=\bigl\||f|\big|_{\Sigma}\bigr\|_{L^{q}(\Sigma)}^{2}\\ &\leq C\bigl\||f|\big|_{\Sigma}\bigr\|_{H^{1/2-\mu/2}(\Sigma)}^{2}\\ &\leq C\bigl\||f|\big|_{\Sigma}\bigr\|_{H^{1/2-\mu^{\prime}/2}(\Sigma)}^{2}\\ &\leq C\bigl\||f|\bigr\|_{H^{1-\mu^{\prime}/2}(\mathbb{R}^{n})}^{2}\\ &\leq\delta\bigl\||f|\bigr\|_{H^{1}(\mathbb{R}^{n})}^{2}+(C_{\delta}-1)\bigl\||f|\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}\\ &\leq\delta\|f\|_{\mathcal{H}_{A,Q}}^{2}+C_{\delta}\|f\|_{L^{2}(\mathbb{R}^{n})}^{2},\end{split}$ (2.10)